34SP.com System Status History

queeg.34sp.com outage (queeg.34sp.com)

resolved

Description

This matter has now been resolved - apologies once again for the disruption to services during this period.

The cause of this has now been fully identified and this should not re-occurr.



The source of these issues has now been traced and we are in active communication with the upstream provider at this time.

We hope to have all services restored very shortly.



We are aware of the problem with the queeg.34sp.com server. We are looking into this, and hope to have service restored as soon as possible.

Start Time

24th February 2005, at 4:50PM

Resolution Time

24th February 2005, at 6:56PM

Queeg routing problem (queeg.34sp.com)

resolved

Description

It seems the routing problem affecting queeg has happened again on the new IP address. We are changing the main IP address of queeg to 217.163.9.55 with immediate effect.

We are changing DNS records on our nameservers to reflect this change, but this will take up to 24 hours to take effect. We are also shortening the cacheing times on records associated with queeg as a precautionary measure.

If you run your own DNS services you will need to update this address yourself.

Start Time

23rd February 2005, at 7:16AM

Resolution Time

23rd February 2005, at 7:16AM

deepthought.34sp.com reboot

resolved

Description

The deepthought.34sp.com server is curretnly being rebooted.

The server went offline at approx 12.50 GMT and is expected back online shortly.

Deepthought is now back online as of 13.05 GMT.

Start Time

21st February 2005, at 12:55PM

Resolution Time

21st February 2005, at 1:05PM

Queeg Routing Problem. (queeg.34sp.com)

resolved

Description

It has come to our attention that a routing problem has occurred to our Queeg server causing a loss of all services.

This issue is with our connection provider and they are looking into a resolution for this at present.

Apologies for any inconvenience caused.

Start Time

18th February 2005, at 9:47AM

Resolution Time

21st February 2005, at 3:26PM

Ongoing spam.34sp.com issues (spam.34sp.com)

resolved

Description

UPDATE 8.50PM GMT - all mail should now be restored to the new server.

Apologies to all users for any disruption this issue caused - do let us know if you have any ongoing issues regarding this.

If people need the addresses from the webmail addressbook please email us as soon as possible with the full email address and we will attempt to recover these for you so that you can reinsert them.





UPDATE 2.50PM GMT - mail services should now be restored on the new server. New mail should slowly start to arrive. Old mail is now being transferred across to your account. It will be a few more hours until old mail will be restored.


Due to the continued issues on spam.34sp.com we are migrating all data and user settings to a new server.

During this migration, incoming mail will not be received and will be queued at the sending server until mail is being accepted onto the new server.

As such no mail should be lost - but you will be unable to access email while the migration is completed.

We expect to complete this procedure by 3-4PM GMT.

Once complete we hope to see stability regained and performance improved. The new server is significantly more powerful than its predecessor.

Start Time

15th February 2005, at 11:49AM

Resolution Time

16th February 2005, at 8:54PM

spam.34sp.com (spam.34sp.com)

resolved

Description

We are aware of an issue on this server and there is currently an engineer enroute to the datacenter to look further into this issue.

This issue has now been corrected.

Start Time

12th February 2005, at 9:04AM

Resolution Time

12th February 2005, at 11:24AM

Holly interuption in service (holly.34sp.com)

resolved

Description

An engineer is currently on the way to look at this server. We will post more information as soon as we have it.

The holly.34sp.com server has been rebooted and failed to recover. The server became unstable after a sustained attack against a user site. We will be looking to mitigate this after the server is back online.

We have found a disk inconsistency on the server and are now running the equivalent of scandisk on the server.

Holly is now available again. We are keeping a close eye on the situation.

We apologise for any inconvenience caused this has caused.

Start Time

11th February 2005, at 3:22PM

Resolution Time

11th February 2005, at 5:41PM

lothos.34sp.com upgrade (lothos.34sp.com)

resolved

Description

Lothos is now undergoing its upgrade, so is unavailable at this time.

Lothos' upgrade is now complete.

Start Time

11th February 2005, at 10:00AM

Resolution Time

11th February 2005, at 11:07AM

kitt.34sp.com upgrade (standard)

resolved

Description

Kitt is now undergoing its upgrade, so is unavailable at this time

Kitt's upgrade is now complete.

Start Time

10th February 2005, at 10:00AM

Resolution Time

10th February 2005, at 11:26AM

gort.34sp.com upgrade (gort.34sp.com)

resolved

Description

Gort is now undergoing its upgrade, so is unavailable at this time.

Gort's upgrade has failed, so it is currently running Plesk 6, as before. We will reschedule this upgrade for next Friday. All gort users will receive a mail to their registered e-mail address to confirm this.

Start Time

10th February 2005, at 10:00AM

Resolution Time

10th February 2005, at 11:48AM